# RuleR: Improving LLM Controllability by Rule-based Data Recycling

### v1

@@ Abstract
Large language models (LLMs) still lack delicate controllability over their responses, which is critical to enhancing their performance and the user experience. However, curating supervised fine-tuning (SFT) datasets to improve LLM controllability usually relies on human experts or proprietary LLMs, which requires additional costs. To bridge this gap, we propose Rule-based Data Recycling (RuleR), a data augmentation method incorporating multiple constraints into the original data samples according to predefined rules, which creates new training tasks to consolidate the controllability of LLMs. Instead of creating new data from scratch, RuleR “recycles” existing data by simply applying rule-based edits to their responses and appending the rule-instructions in their original instructions. Experimental results demonstrate RuleR’s effectiveness in improving LLM controllability while maintaining general instruction-following capabilities. The code will be released on https://github.com/MingLiiii/RuleR.
@@ 1 Introduction
Figure 1: Comparing existing methods (top) and our RuleR (bottom) for enhancing LLM controllability. Most existing methods rely on extra-human/model supervision to generate or edit instructions and responses, neglecting the remaining potential of the original data. On the contrary, RuleR demonstrates that simple rule-based (human/model-free) editing of existing data can greatly improve LLM controllability.
@@
Despite the remarkable advancement of the current LLMs and the continuous efforts to build high-quality supervised fine-tuning (SFT) datasets Brown et al. (2020); OpenAI (2023); Touvron et al. (2023a, b); Jiang et al. (2023), one critical aspect is how these LLMs better interact with human responses to maximize their utility and effectiveness for end-users Liu et al. (2024). According to the systematic investigation from Liu et al. (2024), it is essential for LLMs to constrain their outputs to follow specific user-defined formats or standards. Without constraints on responses, it is hard to directly put the free-formed responses into practical usage. This is also verified in the area of LLM Agents Li et al. (2024f); Chen et al. (2023, 2024b); Zhang et al. (2024), in which a predefined format is always necessary.
@@
However, existing methods for building SFT datasets focus mainly on general instructions Wei et al. (2022); Wang et al. (2022); Taori et al. (2023); Xu et al. (2023); Zhou et al. (2023a); Li et al. (2023a); Zhang et al. (2023); Xu et al. (2024), neglecting the potential user-specific constraints and thus lacking delicate controllability of the response on user-desired lengths or formats Chen et al. (2024a); Xia et al. (2024). One potential method to enhance existing SFT data with multi-constraint controllability is to rewrite or modify instructions and responses by expert humans/models Xu et al. (2023); Li et al. (2023a, 2024b, 2024a); He et al. (2024) as shown in Figure 1 (top). However, these methods directly discard or rewrite the original data. Their data curation is costly and inefficient, requiring careful editing by human experts or proprietary LLMs. This raises the question: Can we “recycle” existing SFT datasets for controllability tuning by enforcing various types of constraints to them without human/model editing?
@@
Drawing inspiration from recent work like Mosaic-IT Li et al. (2024c), which develops a human/model-free method augmenting SFT data, and IFEval Zhou et al. (2023b), which utilizes verifiable constraints to evaluate LLMs’ controllability, we propose Rule-based Data Recycling (RuleR), which automatically edits existing SFT data for improving LLM controllability. The key insight of RuleR is to replace human/model editing with various predefined rules of constraining patterns. Specifically, we manually inspect and construct several verifiable constraint rules, in which the original instructions are appended with various constraints, and the responses are optionally modified to fulfill these constraints based on their original characteristics. Our predefined rules cover a wide range of diverse constraints ranging from high-level constraints, e.g., controlling the word frequency in the response, to lower-level constraints, e.g., setting specific wrapping formats [^1] The complete list of rules and descriptions can be found in the Appendix C. . Each rule contains (1) a set of constraint templates that serve as additional rule-instructions, on constraints the LLM should follow, and (2) specified operations that alternately edit the instruction and response to reach an alignment between them (i.e., all the constraints in the instruction hold in the response). For each sample from the original dataset, we randomly draw several rules to be applied in the editing. This produces an augmented sample with the constraints enforced that can be used for controllability tuning.
An example is shown in Figure 1 (bottom). The original instruction is “What type of fruit would be a good addition to a fruit salad?”, and the corresponding response is “Grapes, kiwis, oranges, apples.”. If the constraint “Num of Nouns” is sampled and implemented on this sample, the augmented instruction becomes “What type of fruit would be a good addition to a fruit salad? Ensure the response has more than 3 nouns.”, and the original response remains unchanged on this type of rule. In this example, the constraint (number of nouns) is directly decided by the original data sample (4 nouns), thus avoiding the regeneration of response.
@@
Extensive experimental results on the IFEval (Instruction-Following Eval) Zhou et al. (2023b) benchmark on various datasets demonstrate the effectiveness of our human/model-free method in enhancing LLM controllability. Moreover, the instruction-following metrics also show that RuleR still preserves the general instruction-following ability promoted by the original SFT dataset. To the best of our knowledge, RuleR is the first human/model-free data augmentation approach designed to improve LLM controllability in enforcing multiple constraints to LLM-generated response.
2 Method
@@ 2.1 Preliminaries
Given a supervised finetuning dataset $D$, there are $n$ data samples each of which is represented by a tuple $(x_{i},y_{i})$, where $x_{i}$ represents the instruction and $y_{i}$ represents the corresponding response. Let $p_{\theta}(\cdot)$ denote the LLMs to be trained, with parameters $\theta$. In the instruction tuning setting, $p_{\theta}$ is typically fine-tuned by maximizing the following objective on each data $(x_{i},y_{i})$, $y_{i,j}$ represents the $j_{th}$ token of response $y_{i}$ and $l_{i}$ represents the token length:
$\max_{\theta}\sum_{j=1}^{l_{i}}\log p_{\theta}\left(y_{i,j}|x_{i}\right),$ (1)
2.2 Rule-based Data Recycling
The basis for augmenting existing SFT datasets is to reconstruct instructions and responses. Most existing methods utilize human experts or strong teacher models for this process, which is costly, inefficient, and fails to get fully used to the existing data, as represented in Figure 1 (top). Inspired by Mosaic-IT which proposes a human/model-free augmentation method to reconstruct instruction-response pairs, we are motivated to design a rule-based method for controllability augmentation.
@@ 2.2 Rule-based Data Recycling
Table 1: The comparisons on the Instruction Following Eval, Pair-wise Comparison Winning Score, and the Open LLM Leaderboard, on 3 different base models and several different instruction tuning datasets. The baselines are models trained with the original dataset. Prompt and Inst represent Prompt-level and Instruction-level accuracy; S and L represent Strict and Loose versions. A, H, M, and T represent ARC, HellaSwag, MMLU, and TruthfulQA.
With this goal, the first challenge is the reformulation of instructions. Due to the lack of human experts or LLMs to generate the corresponding responses for any given instructions, it is not feasible to add arbitrary constraints to the original instructions. To solve this problem, we propose to only incorporate constraints that are compatible with the original data sample. Specifically, we propose to enhance LLM controllability by Rule-based Data Recycling (RuleR), in which the rules and constraints used to reformulate instruction are determined by the original data, ensuring the alignment between instruction and response. Motivated by IFEval, we manually inspect and propose several verifiable rules in which the responses are not required for regeneration by LLMs or human experts, but only need optional slight adjustments. The complete list of rules is listed in Table 3 ranging from high-level constraints like term frequency to lower-level constraints like specific wrapping formats.
During implementation, we instantiate each predefined rule as a tuple , $(\mathbf{S}_{k},f_{k},g_{k})$, where $\mathbf{S}_{k}$ represents the set of manually curated rule-instruction templates for the $k^{th}$ rule, $f_{k}$ and $g_{k}$ represent the function to reformulate instructions and optionally modify responses respectively, as shown in Figure 1(bottom). For each data sample $(x_{i},y_{i})$, the augmented instruction $x_{i,aug}$ will be reformulated according to the characteristics of the original data and the corresponding template sets:
$x_{i,aug}=f_{k}(x_{i},y_{i},\mathbf{S}_{k})$
where specifically, a random rule-instruction template will be sampled and filled with the corresponding characteristics as the additional constraint to the original instruction. Then the rule-instruction is concatenated with the original instruction becoming the augmented instruction. Similarly, the augmented response $y_{i,aug}$ will be optionally modified based on the selected rule $k$:
$y_{i,aug}=g_{k}(x_{i},y_{i},\mathbf{S}_{k}),$
where the function $g_{k}$ varies dramatically for different rules: For some of the rules, no modification needs to be done to the original response, while for others, some arbitrary part of the response needs to be replaced with the desired user format.
3 Experiments
3.1 Experimental Setup
@@
We conduct experiments on several datasets including the Alpaca Taori et al. (2023), Alpaca-GPT4 Peng et al. (2023), WizardLM Xu et al. (2023), Recycled WizardLM Li et al. (2023a), DEITA Liu et al. (2023), and on several base LLMs including Llama2-7B, Llama2-13B Touvron et al. (2023b), and Mistral-7B Jiang et al. (2023), introduction on datasets can be found in the Appendix A. We utilize the prompt and code base from Vicuna Chiang et al. (2023a) and flash attention Dao et al. (2022). The Adam optimizer Kingma and Ba (2017) is utilized with the batch size of $128$ and with the max token length of $2048$. For training on Llama2-7B and Llama2-13B Touvron et al. (2023b), the maximum learning rate is set to $2\times 10^{-5}$ with a warmup rate of $0.03$ for $3$ epochs. For training on Mistral-7B Jiang et al. (2023), the maximum learning rate is set to $1\times 10^{-5}$ with a warmup rate of $0.1$ for $2$ epochs. When utilizing our method, we run the augmentation process $3$/$2$ times to simulate the epochs of training. These augmented data are then mixed together and used for training $1$ epoch. All other configurations are kept the same as the baselines.
3.2 Experimental Results

### v3

@@ Abstract
Despite the remarkable advancement of Large language models (LLMs), they still lack delicate controllability under sophisticated constraints, which is critical to enhancing their response quality and the user experience. While conditional supervised fine-tuning (SFT) can potentially improve LLM controllability, curating new SFT data to fulfill the constraints usually relies on human experts or proprietary LLMs, which is time-consuming and expensive. To bridge this gap, we propose Rule-based Data Recycling (RuleR), a human/LLM-free data augmentation method incorporating multiple constraints into the original SFT data. Instead of creating new responses from scratch, RuleR integrates linguistic or formatting rules into the original instructions and modifies the responses to fulfill the rule-defined constraints. Training on the “recycled” data consolidates LLM capability to generate constrained outputs. Extensive experiments demonstrate RuleR’s effectiveness in improving LLM controllability while maintaining general instruction-following performance. RuleR’s code is released on https://github.com/tianyi-lab/RuleR.
@@ 1 Introduction
Figure 1: Comparing widely-used data generation strategy (top) and RuleR (bottom) enhancing LLM controllability. Most existing methods rely on human/model supervision or reflection to generate new instructions and responses. However, discarding existing data is a waste of efforts taken to curate them by human experts or expert LLMs. On the contrary, our RuleR demonstrates that simple rule-based (human/model-free) editing of existing data can greatly improve LLM controllability.
Figure 2: Examples of the data recycling workflows in Rule-based Data Recycling (RuleR). (a), (b) and (c) select different predefined rules to modify the original data to fulfill constraints on the complexity or format of the response. The differences in new responses are highlighted in red, the example in (a) has already satisfied the appended constraint, thus the response is kept unchanged.
@@
Despite the remarkable advancement of the current Large language models (LLMs) and the continuous efforts to build high-quality supervised fine-tuning (SFT) datasets Brown et al. (2020); OpenAI (2023); Touvron et al. (2023a, b); Jiang et al. (2023), one critical challenge is to generate responses better interacting with humans, with the utility and effectiveness maximized for end-users Liu et al. (2024). According to the systematic investigation from Liu et al. (2024), it is essential for LLMs to constrain their outputs to follow user-specified formats or characteristics. In various practical applications, without any constraint or format being enforced, free-formed responses are not legal or directly applicable. It has also been verified on LLM Agents Li et al. (2024f); Chen et al. (2023, 2024b); Zhang et al. (2024) that enforcing predefined formats is necessary for various tasks.
@@
However, existing SFT datasets are mainly composed of general instructions without user-specified constraints Wei et al. (2022); Wang et al. (2022); Taori et al. (2023); Xu et al. (2023); Zhou et al. (2023a); Li et al. (2023a); Zhang et al. (2023); Xu et al. (2024) and thus result in models lacking delicate controllability of the lengths and format of responses Chen et al. (2024a); Xia et al. (2024). To enhance the utility of existing SFT data in improving the controllability of LLMs, a potential method is to rewrite or modify instructions and responses by experts such as humans/LLMs Xu et al. (2023); Li et al. (2023a, 2024b, 2024a); He et al. (2024) in order to make them fulfilling multiple constraints, as shown in Figure 1 (top). However, the curation of new data is not only costly and inefficient, requiring careful editing by human experts or proprietary LLMs, but also represents a waste of previous efforts: It is impractical to discard all existing data and create brand new data every time we need to add more constraints to the instructions. Hence, we raise the question: Can we “recycle” existing SFT data without human/LLM editing and enforce various types of constraints in order to improve LLM controllability?
Drawing inspiration from IFEval Zhou et al. (2023b), which utilizes verifiable constraints to evaluate LLMs’ controllability, and the human/model-free data augmentation in Mosaic-IT Li et al. (2024c), we propose Rule-based Data Recycling (RuleR), which automatically “re-engineering” existing SFT data for improving LLM controllability. As illustrated in Figure 1 (bottom), the key insight of RuleR is to automatically build constraint-augmented SFT dataset at no cost of human/LLM, by applying predefined rules to enforce constraints in the instructions and responses of the original dataset. Specifically, we manually inspect and construct a diverse set of rules as constraints, which specify the linguistic or formatting constraints on different parts of the response. To modify the original dataset, these rules are appended to the instructions, while the corresponding responses are modified if necessary to fulfill the constraints of these rules.
@@
Our predefined rules cover a wide range of diverse constraints generalizable to many application scenarios, ranging from high-level constraints, e.g., controlling the word frequency in the response, to lower-level constraints, e.g., setting specific wrapping formats of some keywords. Each rule is composed of (1) multiple templates to produce additional instructions enforcing the constraints, and (2) a piece of code [^1] https://github.com/tianyi-lab/RuleR that alternately edits the original instruction and response in order to make the edited response fulfill all the constraints appended to the instruction. For each sample from the original dataset, we randomly draw several rules to be applied to the editing. This produces an augmented sample with the constraints enforced so it can be used for controllability tuning. The complete list of rules and descriptions can be found in Appendix C and Table 4.
Illustrative examples are provided in Figure 2, which showcase (a) a rule that constrains the number of letters of the response; (b) a rule that specifies the case of specific words (if presenting in the response); and (c) a rule that specifies the wrapping format of specific words. To ensure the consistency between the input constraints and the output response, we modify both the instruction and response based on the characteristics of the original data sample. For each sample, we only sample from the rules applicable to the original response, hence avoiding the potential incorrectness of the edited responses.
@@
Extensive experimental results on the IFEval (Instruction-Following Eval) Zhou et al. (2023b) benchmark on various base models and datasets demonstrate the effectiveness of RuleR in enhancing LLM controllability without extra help from human/model. On the other hand, RuleR still preserves the general instruction-following ability promoted by the original SFT dataset. This is demonstrated by the instruction-following metrics (Pair-wise comparison and Open LLM Leaderboard). To the best of our knowledge, RuleR is the first human/model-free data augmentation and recycling approach designed to improve LLM controllability under multiple constraints.
2 Methodology
@@ 2.1 Preliminaries
Given a supervised finetuning dataset $D$, there are $n$ data samples, each represented by a tuple $(x_{i},y_{i})$, where $x_{i}$ represents the instruction and $y_{i}$ represents the corresponding response. Let $p_{\theta}(\cdot)$ denote the LLM with parameters $\theta$ to be trained. In the instruction tuning setting, $p_{\theta}$ is typically fine-tuned by maximizing the following objective on all the $N$ samples as $(x_{i},y_{i})$, in which $y_{i,j}$ represents the $j_{th}$ token of response $y_{i}$, $y_{i,<j}$ represents the tokens before $y_{i,j}$, and $l_{i}$ represents the token length of $y_{i}$:
$\max_{\theta}\sum_{i=1}^{N}\sum_{j=1}^{l_{i}}\log p_{\theta}\left(y_{i,j}|x_{i},y_{i,<j}\right),$ (1)
[Instruction]
Ensure the 2nd paragraph is enclosed in "-| |-" in the response.
Write a story of no more than 150 words about a knight who defends their kingdom from a dragon.
Make sure the 2nd sentence is enclosed within "\$ \$", the 4th sentence is enclosed within "\$\$\$ \$", the 6th sentence is enclosed within "## #", the 8th sentence is enclosed within "*** *" in the response.
[Response]
In the kingdom of Arvania, there lived Sir Cedric, a valiant knight, sworn to defend the land against any threat. \$One day, the kingdom came under attack from a fire-breathing dragon, terrorizing the people and burning homes.\$ The king called on Sir Cedric, charging him to defeat the dragon and restore peace to the land.
\$\$\$-|Wielding his sword and shield, Sir Cedric rode out to face the dragon, fearlessly engaging it in combat.\$ The battle was fierce, with the dragon’s flames clashing against the knight’s steel. ##But Sir Cedric was determined and unrelenting, his bravery and skill unmatched.|- #
Finally, with a mighty blow, Sir Cedric struck the dragon down, ending its reign of terror. ***The people of Arvania rejoiced, celebrating the bravery of their hero.* And Sir Cedric returned in triumph, hailed as the dragon-slaying knight.
[Instruction]
Ensure the 1st sentence is in uppercase, the 2nd sentence is in uppercase, the 3rd sentence is in uppercase in the response.
Keep the word "way" bounded by "@ @" in the response if it appears.
Describe an example of a time you used to influence in a positive way
Keep the 2nd sentence wrapped in "[[ ]]", the 4th sentence wrapped in "*** ***", and the 5th sentence wrapped in "# #" in the response.
[Response]
I RECENTLY HAD A TEAM PROJECT AT WORK WHERE I HAD TO INFLUENCE MY TEAM MEMBERS TO COMPLETE THE TASK IN THE MOST EFFICIENT @WAY@ POSSIBLE. [[I IDENTIFIED AREAS WHERE OUR TEAM COULD IMPROVE, AND I CONSTANTLY COMMUNICATED AND ENCOURAGED MY TEAMMATES TO STAY FOCUSED AND ON TRACK.]] I ALSO SOUGHT FEEDBACK ON HOW WE COULD IMPROVE THE PROCESS. ***Through my efforts, I managed to motivate my team to complete the project ahead of schedule while also maintaining high-quality work.*** #This helped to boost morale and foster collaboration between our team members.# Everyone was very satisfied with the resulting outcome.
Figure 3: Examples with multiple rules selected and implemented. The randomly generated rule-instructions are colored in violet. The upper example is augmented by $2$ different rules (Paragraph Wrapping, and Sentence Wrapping); the bottom example is augmented by $3$ different rules (Sentence Case, Keyword Wrapping, and Sentence Wrapping). The format differences in new responses are highlighted in red.
2.2 Rule-based Data Recycling (RuleR)
2.2.1 Rule Construction
While most existing methods still require human experts or strong teacher LLMs to generate new data (Figure 1 (top)), we aim at “re-engineering” instructions and responses from existing SFT datasets for building controllability-focused datasets, without expensive and time-consuming supervision from humans or LLMs. In the following, we introduce a rule-based approach “Rule-based Data Recycling (RuleR)” to create high-quality data augmentations for improving LLM controllability.
RuleR reformulates the original instructions and responses by applying rule-based edits according to pre-defined constraints. However, not every constraint can be applied to each sample without humans or LLMs. Hence, we only incorporate constraints compatible with the original sample and those can be implemented with simple rectifications like regular expressions. In RuleR, the rules and constraints are elaborately designed and selected to specify or change certain characteristics of the original data, resulting in consistency between each constraint-augmented instruction and the reformulated response.
Specifically, we focus on the characteristics of responses that can be defined by rules, e.g., by checking the 220 distinct linguistic features in the LFTK package Lee and Lee (2023). In addition, we collect constraints from existing works and widely used instruction-tuning datasets. These diverse characteristics include punctuation-, word-, sentence-, paragraph-level occurrences, and frequencies. We construct rules constraining or specifying the characteristics of the original response. In conjunction with the rules containing these characteristics, we also create rules specifying the format of responses to improve LLM’s format-following capability.
To ensure that the rules selected or sampled for each sample are applicable and legal, we apply the following additional protocols: (1) The rules need to be applicable to the original sample. For instance, the rule “Generating a title before giving the response” is not applicable as we can not generate a title without the help of humans or other additional models. (2) The rules should not remove the content of the original response. For instance, the rule “Ensure the word xxx is not shown in the response” is not applicable since we can not directly remove this word from the response as it is part of the original response. (3) The rules should be compatible with the original sample. If the rule is “Ensure there are more than N sentences in the response”, then it can not be applied to samples whose responses have $<N$ sentences. The complete list of rules is provided in Table 4, which covers both high-level constraints such as the term frequency, and lower-level constraints such as specific wrapping formats.
2.2.2 Rule Implementation
To implement each rule as a program editing instructions and responses, we formulate each pre-defined rule as a tuple, $(\mathbf{S}_{k},f_{k},g_{k})$, where $\mathbf{S}_{k}$ represents the set of manually curated templates for creating instructions of the $k^{th}$ rule, while $f_{k}$ and $g_{k}$ are the corresponding functions to reformulate instructions and responses (if necessary), respectively.
Specifically, the function $f_{k}$ selects an appropriate template of a rule for a given data sample and augments the original instruction with an instruction generated by the template. It first selects a subset of rules applicable to the characteristics of the response (e.g., presence of keywords, number of sentences, etc.). Then, we randomly draw one rule out of the subset and create a formatted instruction of the rule from the template. Such rule instruction is appended to the original instruction.
Specifically, for each data sample $(x_{i},y_{i})$, the augmented instruction $x_{i,aug}$ is reformulated according to the characteristics of the original sample and the corresponding template sets, i.e.,
$x_{i,aug}=f_{k}(x_{i},y_{i},\mathbf{S}_{k}),$ (2)
The function $g_{k}$ is designed to modify the response to be consistent with the augmented instruction (after applying function $f_{k}$), i.e., with the rule applied. Applying $g_{k}$ either preserves the original response or revises it. Some rules do not require editing of responses, e.g., keyword appearance, the number of nouns, etc. For rules defining the case or format of certain parts in the response, modifications are needed. The detailed descriptions of each predefined rule can be found in Appendix C. Specifically, the augmented response $y_{i,aug}$ will be optionally modified based on the selected rule $k$:
$y_{i,aug}=g_{k}(x_{i},y_{i},\mathbf{S}_{k}).$ (3)
With the augmented sample $(x_{i,aug},y_{i,aug})$, the training objective becomes:
$\max_{\theta}\sum_{i=1}^{N}\sum_{j=1}^{l_{i,aug}}\log p_{\theta}\left(y_{i,aug,j}|x_{i,aug},y_{i,aug,<j}\right),$ (4)
where $y_{i,aug,j}$ represents the $j$th token of response $y_{i,aug}$ and $l_{i,aug}$ represents its token length.
2.2.3 Multi-rule Implementation
§ 2.2.2 illustrates how RuleR applies one rule to edit each data sample. To create more complex, diverse, and challenging samples, we can extend it to multiple rules randomly drawn from the feasible set of rules. We provide examples of multi-rule augmentation in Figure 3. It shows that the complexity of the augmented data grows dramatically, forcing the LLMs to learn to follow multiple constraints.
@@ 2.2.3 Multi-rule Implementation
Table 1: Main Results. Evaluation on the Instruction Following Eval, Pair-wise Comparison Winning Score, and the Open LLM Leaderboard. We compare RuleR with Baseline for finetuning three base models on several different instruction tuning datasets. Baseline – models trained with the original dataset; Ours – models trained with RuleR-recycled datasets; Prompt – Prompt-level accuracy; Inst – Instruction-level accuracy; S and L represent Strict and Loose versions. A, H, M, and T denote ARC, HellaSwag, MMLU, and TruthfulQA.
3 Experimental Setup
3.1 Implementation Details
We utilize the prompt and code base from Vicuna Chiang et al. (2023a) and flash attention Dao et al. (2022) for all our experiments.
@@ 3.1 Implementation Details
The Adam optimizer Kingma and Ba (2017) is utilized with the batch size of $128$ and with the max token length of $2048$. For training on Llama2-7B and Llama2-13B Touvron et al. (2023b), the maximum learning rate is set to $2\times 10^{-5}$ with a warmup rate of $0.03$ for $3$ epochs. For training on Mistral-7B Jiang et al. (2023), the maximum learning rate is set to $1\times 10^{-5}$ with a warmup rate of $0.1$ for $2$ epochs. When utilizing our method, we run the augmentation process $3$/$2$ times to simulate the epochs of training. These augmented data are then mixed together and used for training $1$ epoch. All other configurations are kept the same as the baselines.
3.2 Training Datasets
We utilize 5 SFT datasets to evaluate the effectiveness of our method: Alpaca dataset Taori et al. (2023), Alpaca-GPT4 dataset Peng et al. (2023), WizardLM dataset Xu et al. (2023), Recycled WizardLM Dataset Li et al. (2023a), DEITA dataset Liu et al. (2023). A detailed illustration of datasets can be found in the Appendix A.
3.3 Evaluation Metrics
We employ 3 commonly accepted metrics for the evaluation, including IFEval (Instruction-Following Eval), Pair-wise Comparison, and Open LLM Leaderboard.
IFEval Zhou et al. (2023b) is the primary evaluation metric employed in our study due to its compatibility with our motivation. It focuses on evaluating how LLMs follow various additional constraints, such as specifying a word count or requiring the inclusion of certain keywords a specified number of times. To avoid the utilization of LLMs during evaluation, it proposes 25 distinct types of verifiable instructions. There are 541 prompts in total and each of them incorporates one or more of these verifiable instructions, ensuring the comprehensiveness of the evaluation. IFEval serves as a great inspiration for our method, and there exist semantical overlappings between their verifiable instructions and our rules. However, the specific prompts used in IFEval are kept unknown in the construction of our rule templates, avoiding potential template leakage. Moreover, IFEval only needs to verify the responses, while our method needs to modify the responses for the training, which pushes this process a step further. Consequently, this benchmark not only facilitates a comprehensive comparison but also provides valuable insights that align with our purpose.
Pair-wise Comparison involves evaluating responses from LLMs like GPT-4, especially in open-domain contexts. This method has shown a notable alignment with human assessments, providing a credible evaluative foundation Zheng et al. (2024); Li et al. (2023b); Sottana et al. (2023). We utilize test instruction sets from WizardLM Xu et al. (2023) and Vicuna Chiang et al. (2023b), comprising 218 and 80 diverse human-curated instructions, respectively. Following the framework by Li et al. (2024e, d), we prompt judging LLM to rate responses on a scale from 1 to 10 across multiple dimensions such as relevance and accuracy. To mitigate positional bias Ko et al. (2020); Wang et al. (2023a), responses are presented in two distinct sequences. Evaluation outcomes are categorized into the ’win-tie-lose’ system for each instruction. One model’s response can be noted as Win/Lose only if it has a Higher/Lower score in both directions. The specific evaluation prompt is presented in the Appendix B.
4 Experimental Results
4.1 Main Results